A variable y decays exponentially over time t from an initial value y_0 toward zero according to y=y_0\exp(-t/T), where T represents the time constant for the decay. At the moment when t=T what is the relationship between y and y_0?

y = _________ y_0

What is the relationship between y and y_0 when t= 6 T?

y = __________ y_0

How many time constants will have elapsed when y=y_0/ 9?

t = ___________ T
Answer #1

at t = T

y = yo exp (- T/T) = e-1 yo

at t = 6T

y = yo exp (- 6T/T) = e-6 yo

y = yo exp (- t/T)

yo /9 = yo exp (- t/T)

- 2.197 = -t/T

t = 2.197 T